# The Neptunian ridge as a natural outcome of high-eccentricity tidal migration

### 17 Apr 2026 v1

@@ 1 Introduction
Soon after the first exoplanet discoveries, a deficit of Neptunian planets at very short orbital periods was identified (Mazeh et al. 2005; Lecavelier Des Etangs 2007), a feature now commonly referred to as the Neptunian desert. This deficit reflects a genuine scarcity of planets rather than an observational bias (Mazeh et al. 2016). Recent occurrence-rate analyses have reinforced this picture by showing that the desert remains an extremely underpopulated region once survey incompleteness is taken into account (Castro-González et al. 2024a; Cui et al. 2026). However, its origin remains an open problem (e.g. Matsakos & Königl 2016; Bourrier et al. 2018; Owen & Lai 2018).
Atmospheric escape can erode gas-rich planets at short orbital separations and has been invoked as a plausible contributor to the desert (Owen & Lai 2018). In parallel, high-eccentricity tidal migration (HEM; e.g. Ford & Rasio 2008; Correia et al. 2011) has been widely discussed as a mechanism shaping the desert. If outer perturbers drive planets to sufficiently small periastrons, tidal dissipation can lead either to disruption or to rapid circularization onto compact final orbits, with characteristic periods comparable to those of the desert edge (e.g. Liu et al. 2013; Mazeh et al. 2016; Matsakos & Königl 2016). Today, it is increasingly thought that an interplay between these mechanisms shapes the properties of close-in Neptunes (e.g. Bourrier et al. 2018; Owen & Lai 2018), with contributions from tidal inflation and orbital decay (e.g. Hallatt & Millholland 2026a, b).
Recently, the observational picture of the “exo-Neptunian landscape” has sharpened. Using occurrence-rate calculations to correct for survey biases, Castro-González et al. (2024a) redefined the boundaries of the Neptunian desert and showed that it is bordered by an overdensity of planets concentrated at $P_{\rm orb}\simeq 3.2$-$5.7$ d, separating the desert from the more sparsely populated, longer-period Neptunian savanna (Bourrier et al. 2023). This feature, referred to as the Neptunian ridge, resembles the hot-Jupiter pile-up, suggesting that the closest-in Neptune- to Jupiter-size planets may share related evolutionary pathways (Castro-González et al. 2024a). The origin of the pile-up remains debated, but HEM is widely considered to play a key role (see Dawson2018, and references therein). Recent observational evidence points to a similarly important contribution in the ridge (e.g. Correia et al. 2020; Bourrier et al. 2023, 2025; Doyle et al. 2025; Vissapragada & Behmard 2025; Handley et al. 2025; Yee et al. 2025; Handley & Batygin 2026).
@@ 2 Tidal survival and circularization
where $\eta$ is a dimensionless encounter parameter that depends on the planetary internal structure and on the hydrodynamics of the tidal interaction. This condition translates into a minimum circularized orbital period after tidal dissipation. Because the limit depends on the planet’s bulk density, it can be mapped into the period-radius plane through a mass-radius relation, defining a tidal survival boundary $P_{\rm tide}(R_{\rm p})$ for planets delivered through HEM (Matsakos & Königl 2016).
Survival alone does not guarantee compact post-HEM orbits. Tidal circularization is highly sensitive to periastron distance, so planets that avoid very close passages may survive but fail to circularize on system timescales (e.g. Fabrycky & Tremaine 2007). Tidal evolution therefore operates most efficiently in a narrow periastron range near the disruption threshold.
@@ 3 A tidal origin for the Neptunian ridge
Figure 1: Confrontation between tidal survival limits and the close-in planet population. (a) Period-radius plane. The solid black curve traces the bias-corrected Neptunian desert boundary and the vertical dashed line marks the ridge limit (Castro-González et al. 2024a). Red and blue dashed curves show tidal survival boundaries $P_{\rm tide}(R_{\rm p};q)$ for the $q=0.84$ and $q=0.16$ mass-radius envelopes, adopting $\eta=4.5$, and the shaded region indicates the survival band. Dotted curves show representative $P_{\rm circ}^{\max}$ tracks with $Q^{\prime}_{\rm p}=10^{5}$ and $M_{\star}=1\,M_{\odot}$ (Owen & Lai 2018): orange and red correspond to $t_{\rm sys}=8$ and 44 Gyr with $a_{0}=1.5$ au along the $q=0.84$ envelope, and blue to $t_{\rm sys}=8$ Gyr with $a_{0}=1.0$ au along the $q=0.16$ envelope. Dash-dotted dark magenta and pink curves show illustrative secular-chaos circularization curves (e.g. Matsakos & Königl 2016), computed using $M(R_{\rm p};q=0.5)$ and perturbing-companion masses of $0.7$ and $0.4\,M_{\rm J}$, respectively. (b) Mass-radius plane with the quantile envelopes $M(R_{\rm p};q)$.
@@ 3.2 A coupled HEM origin of the ridge and desert boundary
The slope of the desert boundary from the sub-Neptune to the super-Neptune/sub-Saturn regime ($1.8R_{\oplus}\lesssim R_{\rm p}\lesssim 6R_{\oplus}$) follows directly from the HEM tidal survival limit when evaluated along the high-density envelope, $P_{\rm tide}(q=0.84)$. This agreement is not enforced by construction, but instead emerges from the tidal survival criterion itself. Adopting a single representative value of the tidal encounter parameter, $\eta=4.5$, which sets the overall period offset, the predicted boundary reproduces the bias-corrected desert edge over this full radius range. We note that $\eta$ should not be interpreted here as a universal disruption constant, but rather as an effective parameter for the onset of strong tidal disruption within the considered framework. In particular, the derived value is tied to the definition of the desert, the empirical mass-radius relation, and the quantile envelopes. Still, the effective $\eta$ inferred here is of the same order of magnitude as those reported in previous works, including $\eta\sim 2.7$ from hydrodynamic calculations for Jupiter-like planets (Guillochon et al. 2011) and $\eta\sim 3.5$ in a related HEM study of the desert (Matsakos & Königl 2016).
@@
Tidal dissipation increases very steeply toward small periastron distances, so efficient circularization is concentrated among trajectories that pass close to the disruption threshold. As a consequence, successful HEM outcomes typically accumulate just exterior to the survival limit (e.g. Fabrycky & Tremaine 2007; Giacalone et al. 2017; Owen & Lai 2018). This concentration is intrinsically stronger for lower-density planets, whose larger disruption radii shift the survival boundary to longer periods. Therefore, the combination of the steep tidal dependence on periastron distance and the finite density dispersion naturally confines surviving HEM outcomes to a narrow period interval that traces the Neptunian ridge. Figure 1 also shows representative maximum circularization periods for the Lidov-Kozai channel and illustrative secular-chaos circularization curves. While these first-order estimates are model-dependent and intended only as qualitative guides, their overall agreement with the empirical desert-ridge structure supports a major role for HEM in setting its geometry.
@@ 4 Discussion
The results presented in this work support a unified tidal interpretation of the Neptunian ridge and the adjacent desert within the HEM scenario, sharpening earlier HEM-based explanations. Matsakos & Königl (2016) showed that tidal-survival and secular-chaos curves can reproduce the basic desert geometry when combined with an empirical mass-radius relation. In a subsequent study, Owen & Lai (2018) showed that post-HEM circularization typically occurs only slightly beyond the disruption boundary, implying a narrow range of final periods. Our key differences are that we confront the HEM framework with the updated, bias-corrected desert-ridge boundaries and explicitly incorporate the observed dispersion in the mass-radius relation, which broadens the survival limit into a finite tidal survival band whose full extent naturally spans the Neptunian ridge. We note that recent models suggest that HEM-driven circularization of ridge Neptunes can often cause runaway tidal inflation, unless the planets are sufficiently metal-rich or tidal dissipation occurs mainly in their outer layers (Hallatt & Millholland 2026b). Atmospheric survival may therefore further limit which HEM products remain in the ridge.
Independent observational trends further support a HEM origin for the Neptunian ridge overdensity. The excited eccentricities, polar orbits, and evidence for atmospheric erosion in several ridge planets are more consistent with a late, dynamically excited origin than with early, smooth disk-driven migration (e.g. Correia et al. 2020; Bourrier et al. 2023, 2025). Host-star properties point in the same direction: ridge and desert Neptunes preferentially orbit metal-rich stars relative to savanna planets (Vissapragada & Behmard 2025), consistent with a higher occurrence of massive companions capable of triggering HEM.
@@ Acknowledgements.
We thank the anonymous referee for a very careful, constructive, and insightful report, which significantly improved the clarity and overall quality of this work. This work has been carried out within the framework of the NCCR PlanetS supported by the Swiss National Science Foundation under grant 51NF40_205606. This project has received funding from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (project Spice Dune, grant agreement No 947634). ACMC acknowledges support from the FCT, Portugal, through the CFisUC project UID/04564/2025, with DOI identifier 10.54499/UID/04564/2025. This work has made use of the NASA Exoplanet Archive (Christiansen et al. 2025) and the Kepler DR25 catalog (Thompson et al. 2018).
@@ A.1 Tidal disruption and minimum circularization period
where the dimensionless parameter $\eta$ encapsulates the effects of planetary internal structure and the hydrodynamics of the encounter (e.g. Guillochon et al. 2011). Following previous related studies, as a first approximation we assume a single value of $\eta$ over the density range of gas-rich planets from sub-Neptunes to Jupiters (e.g. Matsakos & Königl 2016; Owen & Lai 2018).
In HEM scenarios, planets are driven to highly eccentric orbits ($e\rightarrow 1$). If angular momentum is conserved, the final circularised semi-major axis is (Ford & Rasio 2006)
@@ A.2 Lidov-Kozai circularization
Following Owen & Lai (2018), we estimate the maximum orbital separation from which tidal dissipation can circularize a highly eccentric orbit within a system age $t_{\rm sys}$. This estimate is based on the equilibrium-tide formalism in the weak-friction approximation, appropriate for HEM driven by secular perturbations such as Lidov-Kozai cycles (e.g. Fabrycky & Tremaine 2007; Anderson et al. 2016).
@@ A.3 Secular-chaos circularization
Following Matsakos & Königl (2016), we consider the secular-chaos circularization scaling from Wu & Lithwick (2011). The characteristic post-circularization semi-major axis is written as
@@
where $\alpha$ is a dimensionless normalization parameter and $M_{\rm pert}$ is the mass of the perturbing companion. As in Matsakos & Königl (2016), we adopt $\alpha=1/6$, which sets the reference normalization of the scaling.

### 20 Apr 2026 v2

@@ 1 Introduction
Soon after the first exoplanet discoveries, a deficit of Neptunian planets at very short orbital periods was identified (Mazeh et al. 2005; Lecavelier Des Etangs 2007), a feature now commonly referred to as the Neptunian desert. This deficit reflects a genuine scarcity of planets rather than an observational bias (Mazeh et al. 2016). Recent occurrence-rate analyses have reinforced this picture by showing that the desert remains an extremely underpopulated region once survey incompleteness is taken into account (Castro-González et al. 2024a; Cui et al. 2026). However, its origin remains an open problem (e.g. Matsakos and Königl 2016; Bourrier et al. 2018; Owen and Lai 2018).
Atmospheric escape can erode gas-rich planets at short orbital separations and has been invoked as a plausible contributor to the desert (Owen and Lai 2018). In parallel, high-eccentricity tidal migration (HEM; e.g. Ford and Rasio 2008; Correia et al. 2011) has been widely discussed as a mechanism shaping the desert. If outer perturbers drive planets to sufficiently small periastrons, tidal dissipation can lead either to disruption or to rapid circularization onto compact final orbits, with characteristic periods comparable to those of the desert edge (e.g. Liu et al. 2013; Mazeh et al. 2016; Matsakos and Königl 2016). Today, it is increasingly thought that an interplay between these mechanisms shapes the properties of close-in Neptunes (e.g. Bourrier et al. 2018; Owen and Lai 2018), with contributions from tidal inflation and orbital decay (e.g. Hallatt and Millholland 2026a, b).
Recently, the observational picture of the “exo-Neptunian landscape” has sharpened. Using occurrence-rate calculations to correct for survey biases, Castro-González et al. (2024a) redefined the boundaries of the Neptunian desert and showed that it is bordered by an overdensity of planets concentrated at $P_{\rm orb}\simeq 3.2$-$5.7$ d, separating the desert from the more sparsely populated, longer-period Neptunian savanna (Bourrier et al. 2023). This feature, referred to as the Neptunian ridge, resembles the hot-Jupiter pile-up, suggesting that the closest-in Neptune- to Jupiter-size planets may share related evolutionary pathways (Castro-González et al. 2024a). The origin of the pile-up remains debated, but HEM is widely considered to play a key role (see Dawson and Johnson 2018, and references therein). Recent observational evidence points to a similarly important contribution in the ridge (e.g. Correia et al. 2020; Bourrier et al. 2023, 2025; Doyle et al. 2025; Vissapragada and Behmard 2025; Handley et al. 2025; Yee et al. 2025; Handley and Batygin 2026).
@@ 2 Tidal survival and circularization
where $\eta$ is a dimensionless encounter parameter that depends on the planetary internal structure and on the hydrodynamics of the tidal interaction. This condition translates into a minimum circularized orbital period after tidal dissipation. Because the limit depends on the planet’s bulk density, it can be mapped into the period-radius plane through a mass-radius relation, defining a tidal survival boundary $P_{\rm tide}(R_{\rm p})$ for planets delivered through HEM (Matsakos and Königl 2016).
Survival alone does not guarantee compact post-HEM orbits. Tidal circularization is highly sensitive to periastron distance, so planets that avoid very close passages may survive but fail to circularize on system timescales (e.g. Fabrycky and Tremaine 2007). Tidal evolution therefore operates most efficiently in a narrow periastron range near the disruption threshold.
@@ 3 A tidal origin for the Neptunian ridge
Figure 1: Confrontation between tidal survival limits and the close-in planet population. (a) Period-radius plane. The solid black curve traces the bias-corrected Neptunian desert boundary and the vertical dashed line marks the ridge limit (Castro-González et al. 2024a). Red and blue dashed curves show tidal survival boundaries $P_{\rm tide}(R_{\rm p};q)$ for the $q=0.84$ and $q=0.16$ mass-radius envelopes, adopting $\eta=4.5$, and the shaded region indicates the survival band. Dotted curves show representative $P_{\rm circ}^{\max}$ tracks with $Q^{\prime}_{\rm p}=10^{5}$ and $M_{\star}=1\,M_{\odot}$ (Owen and Lai 2018): orange and red correspond to $t_{\rm sys}=8$ and 44 Gyr with $a_{0}=1.5$ au along the $q=0.84$ envelope, and blue to $t_{\rm sys}=8$ Gyr with $a_{0}=1.0$ au along the $q=0.16$ envelope. Dash-dotted dark magenta and pink curves show illustrative secular-chaos circularization curves (e.g. Matsakos and Königl 2016), computed using $M(R_{\rm p};q=0.5)$ and perturbing-companion masses of $0.7$ and $0.4\,M_{\rm J}$, respectively. (b) Mass-radius plane with the quantile envelopes $M(R_{\rm p};q)$.
@@ 3.2 A coupled HEM origin of the ridge and desert boundary
The slope of the desert boundary from the sub-Neptune to the super-Neptune/sub-Saturn regime ($1.8R_{\oplus}\lesssim R_{\rm p}\lesssim 6R_{\oplus}$) follows directly from the HEM tidal survival limit when evaluated along the high-density envelope, $P_{\rm tide}(q=0.84)$. This agreement is not enforced by construction, but instead emerges from the tidal survival criterion itself. Adopting a single representative value of the tidal encounter parameter, $\eta=4.5$, which sets the overall period offset, the predicted boundary reproduces the bias-corrected desert edge over this full radius range. We note that $\eta$ should not be interpreted here as a universal disruption constant, but rather as an effective parameter for the onset of strong tidal disruption within the considered framework. In particular, the derived value is tied to the definition of the desert, the empirical mass-radius relation, and the quantile envelopes. Still, the effective $\eta$ inferred here is of the same order of magnitude as those reported in previous works, including $\eta\sim 2.7$ from hydrodynamic calculations for Jupiter-like planets (Guillochon et al. 2011) and $\eta\sim 3.5$ in a related HEM study of the desert (Matsakos and Königl 2016).
@@
Tidal dissipation increases very steeply toward small periastron distances, so efficient circularization is concentrated among trajectories that pass close to the disruption threshold. As a consequence, successful HEM outcomes typically accumulate just exterior to the survival limit (e.g. Fabrycky and Tremaine 2007; Giacalone et al. 2017; Owen and Lai 2018). This concentration is intrinsically stronger for lower-density planets, whose larger disruption radii shift the survival boundary to longer periods. Therefore, the combination of the steep tidal dependence on periastron distance and the finite density dispersion naturally confines surviving HEM outcomes to a narrow period interval that traces the Neptunian ridge. Figure 1 also shows representative maximum circularization periods for the Lidov-Kozai channel and illustrative secular-chaos circularization curves. While these first-order estimates are model-dependent and intended only as qualitative guides, their overall agreement with the empirical desert-ridge structure supports a major role for HEM in setting its geometry.
@@ 4 Discussion
The results presented in this work support a unified tidal interpretation of the Neptunian ridge and the adjacent desert within the HEM scenario, sharpening earlier HEM-based explanations. Matsakos and Königl (2016) showed that tidal-survival and secular-chaos curves can reproduce the basic desert geometry when combined with an empirical mass-radius relation. In a subsequent study, Owen and Lai (2018) showed that post-HEM circularization typically occurs only slightly beyond the disruption boundary, implying a narrow range of final periods. Our key differences are that we confront the HEM framework with the updated, bias-corrected desert-ridge boundaries and explicitly incorporate the observed dispersion in the mass-radius relation, which broadens the survival limit into a finite tidal survival band whose full extent naturally spans the Neptunian ridge. We note that recent models suggest that HEM-driven circularization of ridge Neptunes can often cause runaway tidal inflation, unless the planets are sufficiently metal-rich or tidal dissipation occurs mainly in their outer layers (Hallatt and Millholland 2026b). Atmospheric survival may therefore further limit which HEM products remain in the ridge.
Independent observational trends further support a HEM origin for the Neptunian ridge overdensity. The excited eccentricities, polar orbits, and evidence for atmospheric erosion in several ridge planets are more consistent with a late, dynamically excited origin than with early, smooth disk-driven migration (e.g. Correia et al. 2020; Bourrier et al. 2023, 2025). Host-star properties point in the same direction: ridge and desert Neptunes preferentially orbit metal-rich stars relative to savanna planets (Vissapragada and Behmard 2025), consistent with a higher occurrence of massive companions capable of triggering HEM.
@@ Acknowledgements.
We thank the anonymous referee for a very careful, constructive, and insightful report, which significantly improved the clarity and overall quality of this work. This work has been carried out within the framework of the NCCR PlanetS supported by the Swiss National Science Foundation under grant 51NF40_205606. This project has received funding from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (project Spice Dune, grant agreement No 947634). ACMC acknowledges support from the FCT, Portugal, through the CFisUC project UID/04564/2025, with DOI identifier 10.54499/UID/04564/2025. This work has made use of the NASA Exoplanet Archive (Christiansen2025) and the Kepler DR25 catalog (Thompson2018).
@@ A.1 Tidal disruption and minimum circularization period
where the dimensionless parameter $\eta$ encapsulates the effects of planetary internal structure and the hydrodynamics of the encounter (e.g. Guillochon et al. 2011). Following previous related studies, as a first approximation we assume a single value of $\eta$ over the density range of gas-rich planets from sub-Neptunes to Jupiters (e.g. Matsakos and Königl 2016; Owen and Lai 2018).
In HEM scenarios, planets are driven to highly eccentric orbits ($e\rightarrow 1$). If angular momentum is conserved, the final circularised semi-major axis is (Ford and Rasio 2006)
@@ A.2 Lidov-Kozai circularization
Following Owen and Lai (2018), we estimate the maximum orbital separation from which tidal dissipation can circularize a highly eccentric orbit within a system age $t_{\rm sys}$. This estimate is based on the equilibrium-tide formalism in the weak-friction approximation, appropriate for HEM driven by secular perturbations such as Lidov-Kozai cycles (e.g. Fabrycky and Tremaine 2007; Anderson et al. 2016).
@@ A.3 Secular-chaos circularization
Following Matsakos and Königl (2016), we consider the secular-chaos circularization scaling from Wu and Lithwick (2011). The characteristic post-circularization semi-major axis is written as
@@
where $\alpha$ is a dimensionless normalization parameter and $M_{\rm pert}$ is the mass of the perturbing companion. As in Matsakos and Königl (2016), we adopt $\alpha=1/6$, which sets the reference normalization of the scaling.